Crossfire Company segments its business into two regions-East and West. The company prepared a contribution format segmented income statement as shown below: Total Company East West Sales Variable expenses Contribution margin Traceable fixed expenses Segment margin Common fixed expenses Net operating income $1,040,000 676,000 364,000 148,000 216,000 76,000 $ 140,000 $ 650,000 442,000 208,000 64,000 $ 144,000 $ 390,000 234,000 156,000 84,000 $ 72,000 Required 1. Compute the companywide break-even point in dollar sales 2. ℃ompute tne break-even point in dollar sales tor the East region 3. Compute the break-even point in dollar sales for the West region 4. Prepare a new segmented income statement based on the break-even dollar sales that you computed in requirements 2 and 3. Use the same format as shown above. What is Crossfires net operating income (loss) in your new segmented income statement? 5. Do you think that Crossfire should allocate its common fixed expenses to the East and West regions when computing the break- even points for each region?
Answer #1

Solution:

Computation of Contribution Margin Ratio
Total Company East Region West Region
Sales $10,40,000 $6,50,000 $3,90,000
Contribution Margin $3,64,000 $2,08,000 $1,56,000
Contribution margin Ratio (Contribution margin/ sales) 35% 32% 40%

Solution 1:

Company wide Break even point in dollar sales = Company-wide total fixed cost / Company-wide contribution margin ratio

= ($148,000+ $76,000) / 35% = $640,0000

Solution 2:

Break even point in dollar sales for East Region = Fixed Cost traceable to East Region / Contribution Margin Ratio of East region

= $64000/ 32% = $200,000

Solution 3:

Break even point in dollar sales for West Region = Fixed Cost traceable to West Region / Contribution Margin Ratio of West region

= $84000/ 40% = $210,000

Solution 4:

Segmanted Income Statement - Based on break even points of both regions
Particulars Total Company East West
Sales $4,10,000.00 $2,00,000.00 $2,10,000.00
Variable expenses $2,62,000.00 $1,36,000.00 $1,26,000.00
Contribution margin (32%, 40% of sales) $1,48,000.00 $64,000.00 $84,000.00
Traceable Fixed Expenses $1,48,000.00 $64,000.00 $84,000.00
Segment Margin $0.00 $0.00 $0.00
Common fixed expenses $76,000.00
Net Operating Income (Loss) -$76,000.00

Solution 5:

No, Crossfire should not allocate its common fixed Expenses to the regions when computing the break even point for each region.